OpenStudy (anonymous):

simplify square root of 11 over 121

OpenStudy (solomonzelman):

\[~~~~~~~~\frac{\sqrt{11}}{121}=\frac{\sqrt{11}}{11 \times 11}=\frac{\sqrt{11}}{\sqrt{11} \times \sqrt{11} \times \sqrt{11} \times \sqrt{11}}~~~\] all I can do, but square root of 11 over 121 is really most simplified.

OpenStudy (solomonzelman):

I can say \[\sqrt{11}^{\frac{1}{4}}\]